OpenStudy (anonymous):

f(x)5x^2-8+13 and g(x)=7x-6, what is f(x)+g(x)

OpenStudy (anonymous):

Check if your post correct, pls!

OpenStudy (anonymous):

If f(x)= 5x^2-8x+13 and g(x)=7x+16,what is F(x)+g(x)

OpenStudy (anonymous):

Now, combine them terms with same exponent of x!

OpenStudy (anonymous):

f(x) = 5x^2 - 8x + 13 and g(x) = 7x + 16 we get f(x) + g(x) = 5x^2 - 8x + 13 + 7x + 16 f(x) + g(x) = 5x^2 - x + 29